OpenStudy (anonymous):

Find slope of line tangent to f(x) = 11x^2 - 9x + 3 with x = 2.

OpenStudy (jdoe0001):

get the its derivative, using the power rule, set "x" to 0 on the derivative

OpenStudy (anonymous):

im confused mate

OpenStudy (jdoe0001):

equation of slope of a line tangent = derivative of original equation

OpenStudy (jdoe0001):

woops, I meant, set x = 2 rather

OpenStudy (anonymous):

mate, my brain... ya on the right track

OpenStudy (jdoe0001):

well, have you covered derivatives yet? do you know what a derivative is?

OpenStudy (anonymous):

yeah its the rate of change mate

OpenStudy (jdoe0001):

get the its derivative, using the power rule, set "x" to 2 on the derivative

OpenStudy (anonymous):

will ya guide me through it mate? this review sheets confusin' me

OpenStudy (jdoe0001):

well... haemm do you know the power rule?

OpenStudy (anonymous):

not sure don't think so mate

OpenStudy (jdoe0001):

hmmm are you supposed to use .... something ... to solve this?

OpenStudy (jdoe0001):

other than the power rule that is

OpenStudy (anonymous):

nah its justa made up review by the teacher. i dont get it . no guidance

OpenStudy (jdoe0001):

hmmm review = re view = to view again = re visit meaning it has been covered already it seems

OpenStudy (anonymous):

mate i need help. will ya help guide me?

OpenStudy (jdoe0001):

well.... not sure .... what to use =).. .the limit definition of a derivative... which you likely dunno or the power rule which you dunno either... so dunno what applies to your term

myininaya (myininaya):

Are you suppose to use the definition of derivative, you know the one that involves the word limit?

OpenStudy (anonymous):

implicits

myininaya (myininaya):

What do you mean by implicit? Have you ever seen the following rule: \[\frac{d}{dx} x^n=nx^{n-1} \]

OpenStudy (anonymous):

nah mate

OpenStudy (jdoe0001):

you can always check getting it by using its limit defintion -> http://www.youtube.com/watch?v=WERgBgi4hg4

OpenStudy (anonymous):

confusion hurts mate

OpenStudy (jdoe0001):

if you haven't covered the material, it'd not be proper to try do the exercise

OpenStudy (anonymous):

it'd be nice if i cold see that aanwr worked out so i can study fromit

myininaya (myininaya):

If you don't know power rule then you will have to use the definition of derivative: \[\frac{d}{dx} f(x)=\lim_{h \rightarrow 0}\frac{f(x+h)-f(x)}{h}\]

myininaya (myininaya):

Have you see this above definition?

OpenStudy (jdoe0001):

that video covers it that, it has an example worked out

OpenStudy (anonymous):

it should give mea somethin/somethin answer

OpenStudy (anonymous):

bagajr my hero

OpenStudy (anonymous):

so its only 35?

OpenStudy (anonymous):

Its only 35? thats it?

OpenStudy (anonymous):

Yes, a slope is a number.

OpenStudy (anonymous):

Oh so its not 35 by something? I never knew

myininaya (myininaya):

Please don't just give the answer next time. I really have one learn it and then try their assignment. This student doesn't even know the most basic definition of a derivative. It doesn't take a lot of effort to have seen the definition before. I feel there has not been a lot of effort here honestly.

OpenStudy (anonymous):

myiniaya i am just stumped

OpenStudy (anonymous):

@myininaya Yeah but I don't think ripping the answers away from him and shoving equations and comments that he doesn't understand helps either. You can't force them to learn, you have to make them want to.

OpenStudy (anonymous):

thanks ill study it bagajr thanks

myininaya (myininaya):

It is against the rules to use the site for cheating.

OpenStudy (anonymous):

ya but its justa review sheet its not fora grade mate

OpenStudy (anonymous):

@jeffsours Ask me if you have any specific questions about this. But do you understand what I did up there? When you take a derivative of x^n, just multiply by n (the power) and subtract 1 from the power.\[\Large x^n \rightarrow nx^{n-1}\]\[\Large x^3 \rightarrow 3x^2\]\[\Large x^2 \rightarrow 2x^1=2x\]\[\Large11x^2 \rightarrow 22x^1=22x\]\[\Large9x \rightarrow 9x^0=9*1=9\]Also, derivative of any constant is 0 because derivative is how much something changes and constans don't change so\[\Large3 \rightarrow 0\]\[\Large1 \rightarrow 0\]\[\Large \pi \rightarrow 0\]If you don't understand all of this, please ask because this is the bare basics and you really ned to know it.

OpenStudy (anonymous):

With that, I think i undestand it better. im gonna print your response out as notes

OpenStudy (anonymous):

okay good